# IRCAN: Mitigating Knowledge Conflicts in LLM Generation via Identifying and Reweighting Context-Aware Neurons

### v1

@@ Abstract
It is widely acknowledged that large language models (LLMs) encode a vast reservoir of knowledge after being trained on mass data. Recent studies disclose knowledge conflicts in LLM generation, wherein outdated or incorrect parametric knowledge (i.e., encoded knowledge) contradicts new knowledge provided in the context. To mitigate such knowledge conflicts, we propose a novel framework, IRCAN (Identifying and Reweighting Context-Aware Neurons) to capitalize on neurons that are crucial in processing contextual cues. Specifically, IRCAN first identifies neurons that significantly contribute to context processing, utilizing a context-aware attribution score derived from integrated gradients. Subsequently, the identified context-aware neurons are strengthened via reweighting. In doing so, we steer LLMs to generate context-sensitive outputs with respect to the new knowledge provided in the context. Extensive experiments conducted across a variety of models and tasks demonstrate that IRCAN not only achieves remarkable improvements in handling knowledge conflicts but also offers a scalable, plug-and-play solution that can be integrated seamlessly with existing models.
@@ 1 Introduction
Large language models (LLMs), trained on extensive data, are known for encapsulating a broad spectrum of knowledge. However, due to the rapid evolution of information/knowledge as well as noise in training data, LLMs may possess incorrect or outdated knowledge. To mitigate this issue, in real-world applications, methods like retrieval-augmented generation (RAG) are usually used to integrate latest event knowledge or knowledge from external databases into the context of prompts fed into LLMs. This enables online updates to knowledge and the incorporation of domain-specific information, enhancing the accuracy and reliability of the outputs of LLMs.
@@
We conduct extensive experiments on a diverse array of models from multiple families, including LLaMA [33], Gemma [27] and Amber [23], spanning various parameter scales (2B, 7B, 8B, 13B) and encompassing both pre-trained and instruction-tuned models. To conduct a comprehensive evaluation, we carry out experiments on two types of tasks: completion and multiple choice. Experiment results demonstrate that our method can effectively identify neurons responsible for processing the context within LLMs. Moreover, by enhancing these neurons, LLMs can be guided to remain more faithful to the information provided in the context when facing knowledge conflicts, rather than sticking to its intrinsic knowledge. Additionally, our method can serve as a plug-and-play module, easily integrated with existing approaches. In completion tasks, IRCAN has achieved state-of-the-art performance, with substantial improvements of 129% and 136% in terms of accuracy for LLaMA-2-7B and LLaMA-3-8B respectively. Remarkably, when our method is integrated with CAD [30], a previous strong method for dealing with knowledge conflicts, the performance of LLMs can be further improved. In multiple-choice tasks, IRCAN achieves comparable results to the baseline, and when combined with CAD, our method sets new state-of-the-art results.
@@
  • We conduct experiments to demonstrate that the proposed approach can significantly boost the performance of LLMs on tasks involving knowledge conflicts.
@@ 2 Related Work
To correct outdated or incorrect knowledge in language models, three main lines of strategies have been explored in previous works.
@@
Fine-tuning aims to update the internal knowledge of an existing LLM through further training on additional data, including datasets with the latest information or domain-specific datasets [18, 10, 36, 15]. However, this process requires substantial computational resources and a large amount of training data, as well as significant training time, which can be unaffordable in many cases. More seriously, it may lead to catastrophic forgetting issues.
@@
Model Editing seeks to edit incorrect or undesirable knowledge encoded in pre-trained models. Zhu et al. [41] employ a constrained fine-tuning strategy to fine-tune the model on new data under certain preset constraints. Other works initially identify knowledge-related parameters of the existing pre-trained models and then directly edit the particular knowledge into these parameters [11, 25, 26]. Additionally, efforts have been also explored to store new or correct knowledge in an extra memory, replacing the original predictions with this knowledge during generation [28, 8]. However, these approaches are only applicable to modifying specific knowledge. In contrast, our method is independent of specific knowledge: regardless of the type of knowledge contained in the context, it enhances the LLM’s utilization of this knowledge.
@@ 3 Methodology
We focus on tasks involving context-specific knowledge conflicts. The input of these tasks is formulated as $(c,q)$, where $c$ is the context, and $q$ represents the question in completion tasks or the question combined with a suffix consisting of choices in multiple-choice tasks. We propose a novel method that is dedicated to improving the faithfulness of LLMs to the context to address these tasks. The proposed IRCAN methodology is structured into three phases: Initially, we compute the attribution scores of each neuron to assess its influence on context processing. Subsequently, neurons that are responsive to context, termed “context-aware neurons”, are identified. In the final step, we enhance the influence of these detected neurons through a reweighting process, thereby augmenting their impact on the model’s generation. The framework of IRCAN is shown in Figure 1.
Figure 1: The diagram of IRCAN. When a LLM faces a knowledge conflict between the context and its inherent knowledge, IRCAN first calculates the attribution score for each neuron to measure its contribution to processing context. It then identifies context-aware neurons by taking the intersection of neurons with the highest scores. Finally, the impact of these identified neurons is amplified by increasing their weights, thereby ensuring the model more faithful to the knowledge in the context.
@@ 3.2 Context-Aware Neuron Identification
Based on the calculated neuron attribution scores $\mathrm{Attr}(n_{i}^{l})$, we first retain neurons with scores above the threshold $t$, creating a coarse set of context-aware neurons. Then, for each example, we select the top $z$ neurons with the highest attribution scores as the candidate set. In our experiments, $t$ and $z$ are set to 10% and 20 respectively. Ultimately, we allow each example to vote for the candidate neurons based on their attribution scores, and we select the top $h$ neurons that receive the most votes. These neurons, as the identified context-aware neurons, are enhanced to make LLMs more context-focused.
@@ 4.1 Dataset
For the multiple-choice task, we utilized the COSE_KRE and ECARE_KRE datasets Ying et al. [37], which were derived from the ECQA [1] and e-CARE [9] datasets, respectively. The derivation process involved selecting one of the incorrect answer choices and prompting ChatGPT to generate explanations supporting this incorrect answer. Specifically, the selected incorrect answer was treated as the correct answer, and the explanations generated by ChatGPT were used as the context for the multiple-choice question.
@@ 4.2 Models and Metrics
Although we used accuracy as the primary metric to evaluate our method more comprehensively, we also designed a supplementary metric: stubbornness rate (SR), which measures whether the LLM persistently adheres to its internal memorized knowledge. This metric is defined as the accuracy with which the model’s generation matches the common word of a well-known proverb (for the MemoTrap dataset, i.e., the common ending word; for the COSE_KRE and ECARE_KRE dataset, i.e., the original golden option). A lower stubbornness rate indicates that the LLM exhibits a decreased reliance on the knowledge encapsulated within its internal parameters during the generation process, suggesting a greater propensity to incorporate contextual information to a certain extent.
@@ 4.3 Baselines
To demonstrate the effectiveness of IRCAN, we compared it with the following baselines: Original, which refers to the LLMs without any modification. Context-aware decoding (CAD) [30]: it adjusts the output probabilities of LLMs to emphasize differences when the context is utilized versus when it is absent. Specifically, this approach reduces the weighting of prior knowledge in favor of more pertinent contextual information. Since this method and our approach work in completely different ways: CAD manipulates the model’s final output probabilities, while IRCAN focuses on identifying and enhancing neurons responsible for processing the context, these two methods can be seamlessly combined without any obstacles. In our experiments, we also report the performance of the combined system, which is referred to as IRCAN-CAD in our experiments. Opin [40]: this method argues that transforming factual questions into questions seeking opinions allows people to pay more attention to the context. Therefore, it uses opinion-based prompts to enable the model to generate context-faithful responses. Since the prompts designed by Opin do not apply to the completion task, we used this method as a baseline only on the multiple-choice datasets. To ensure a fair comparison, we adopted the same method for selecting the generated answer as described in §4.2 for Opin.
@@ 4.4 Main Results
We treat the number of identified context-aware neurons $h$3.2), ranging from 1 to 16, and the enhancement strength $\beta$ for these neurons (§3.3), ranging from 2 to 20, as hyperparameters. All datasets were randomly divided into validation and test sets in a 1:1 ratio. We conducted a grid search to identify the optimal hyperparameter configuration that maximized performance on the validation set. Subsequently, we report the test set results obtained using the identified optimal hyperparameter combination in Table 1 and Table 2.
@@
Main results on the MemoTrap dataset are shown in Table 1. Notably, IRCAN significantly outperforms both the Original and the CAD baseline. Particularly, IRCAN achives remarkable improvements of 129% for LLaMA-2-7B and 136% for LLaMA-3-8B in terms of accuracy when compared to the Original baseline. Such substantial performance improvements, achieved through strengthening merely a few or a dozen neurons, indicate that the neurons identified by our method play a pivotal role in processing context.
@@
As presented in Table 2, IRCAN significantly enhances the performance of LLMs in resolving knowledge conflicts in the multiple-choice task. This improvement is observed across the COSE_KRE and ECARE_KRE datasets for the majority of the evaluated LLMs. The multiple-choice task poses a greater challenge compared to the MemoTrap dataset, where the context directly provides the required knowledge for generati-
Figure 3: Model performance with different $\beta$.
Figure 4: Model performance with different $h$.
on. In the multiple-choice task, LLMs are required to interpret and reason based on the implicit knowledge in the context to facilitate generation. Consequently, the improvements in performance observed for the two baseline models and our proposed IRCAN are less pronounced on this dataset than those noted on MemoTrap. Nonetheless, IRCAN still achieves a gain in accuracy and a drop in SR for the majority of LLMs engaged in this task, setting new state-of-the-art results.
@@ 5.1 Effect of Enhancement Strength
The relationship between the accuracy of the LLaMA-2-7B model on the MemoTrap dataset and the enhancement strength $\beta$ is depicted in Figure 3, where the number of enhanced neurons is fixed at 14. For further results on various enhancement strengths $\beta$, please refer to Figure 11 in Appendix C. It can be observed that as the enhancement strength for the context-aware neurons increases, model performance gradually improves, highlighting the pivotal role of the neurons identified by our method. Then, consistent with our intuition, performance begins to decline beyond a certain enhancement strength (7 in this scenario). This decline could be due to excessively high enhancements of certain neurons, leading to uncontrollable outputs or a reduction in model capabilities.
@@ 5.2 Effect of the Number of Context-Aware Neurons
The impact of the number of enhanced neurons $h$ on the performance of LLaMA-2-7B on the MemoTrap dataset is shown in Figure 4. We present results where the enhancement multiplier is fixed at 7, with comprehensive results available in Appendix C (see Figure 12). Similarly, observations reveal that as the number of enhanced neurons increases, the model’s accuracy initially improves but subsequently begins to decline, resonating with the results observed with the enhancement strength.
@@ 5.3 Layer Distribution of Context-Aware Neurons
We also analyzed the overlap of identified context-aware neurons across different prompts. Results (Appendix D) show a high degree of overlap, demonstrating IRCAN consistently identifies these neurons regardless of the prompt.
@@ 5.4 Evaluation of General Abilities of LLMs
To investigate whether up-weighting context-aware neurons impairs the model’s general abilities, we conducted evaluations of IRCAN on six widely-used benchmarks. These benchmarks are used in the widely-recognized Open LLM Leaderboard, [^3] https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard including ARC [5], HellaSwag [38], MMLU [13], Winogrande [29], GSM8K [6], and TruthfulQA [21]. We describe the experimental details in Appendix E. The experimental results, as shown in Table 3, reveal that IRCAN rarely impacts the general ability of the LLMs. Surprisingly, in some cases, it even leads to a slight performance improvement. These results suggest that IRCAN can reliably improve the capability of the LLMs in addressing knowledge conflict tasks while maintaining their excellent general capabilities.

### v2

@@ Abstract
It is widely acknowledged that large language models (LLMs) encode a vast reservoir of knowledge after being trained on mass data. Recent studies disclose knowledge conflicts in LLM generation, wherein outdated or incorrect parametric knowledge (i.e., encoded knowledge) contradicts new knowledge provided in the context. To mitigate such knowledge conflicts, we propose a novel framework, IRCAN (Identifying and Reweighting Context-Aware Neurons) to capitalize on neurons that are crucial in processing contextual cues. Specifically, IRCAN first identifies neurons that significantly contribute to context processing, utilizing a context-aware attribution score derived from integrated gradients. Subsequently, the identified context-aware neurons are strengthened via reweighting. In doing so, we steer LLMs to generate context-sensitive outputs with respect to the new knowledge provided in the context. Extensive experiments conducted across a variety of models and tasks demonstrate that IRCAN not only achieves remarkable improvements in handling knowledge conflicts but also offers a scalable, plug-and-play solution that can be integrated seamlessly with existing models. Our codes are released at https://github.com/danshi777/IRCAN.
@@ 1 Introduction
Large language models (LLMs), trained on extensive data, are known for encapsulating a broad spectrum of knowledge [13, 40, 26, 35]. However, due to the rapid evolution of information/knowledge as well as noise in training data, LLMs may possess incorrect or outdated knowledge. To mitigate this issue, in real-world applications, methods like retrieval-augmented generation (RAG) are usually used to integrate latest event knowledge or knowledge from external databases into the context of prompts fed into LLMs. This enables online updates to knowledge and the incorporation of domain-specific information, enhancing the accuracy and reliability of the outputs of LLMs.
@@
We conduct extensive experiments on a diverse array of models from multiple families, including LLaMA [39], Gemma [31] and Amber [27], spanning various parameter scales (2B, 7B, 8B, 13B) and encompassing both pre-trained and instruction-tuned models. To conduct a comprehensive evaluation, we carry out experiments on two types of tasks: completion and multiple-choice. Experiment results demonstrate that our method can effectively identify neurons responsible for processing the context within LLMs. Moreover, by enhancing these neurons, LLMs can be guided to remain more faithful to the information provided in the context when facing knowledge conflicts, rather than sticking to its intrinsic knowledge. Additionally, our method can serve as a plug-and-play module, easily integrated with existing approaches. In completion tasks, IRCAN has achieved state-of-the-art performance, with substantial improvements of 129% and 136% in terms of accuracy for LLaMA-2-7B and LLaMA-3-8B respectively. Remarkably, when our method is integrated with CAD [36], a previous strong method for dealing with knowledge conflicts, the performance of LLMs can be further improved. In multiple-choice tasks, IRCAN achieves comparable results to the baseline, and when combined with CAD, our method sets new state-of-the-art results.
@@
  • We conduct extensive experiments and experiment results demonstrate that the proposed approach can significantly boost the performance of LLMs on tasks involving knowledge conflicts.
@@ 2 Related Work
To correct outdated or incorrect knowledge in language models, previous studies have explored three main strategies: fine-tuning, model editing and contrastive decoding.
@@
Fine-tuning aims to update the internal knowledge of an existing LLM through further training on additional data, including datasets with the latest information or domain-specific datasets [50, 20, 11, 45, 17]. However, this process requires substantial computational resources and a large amount of training data, as well as significant training time, which can be unaffordable in many cases. More seriously, it may lead to catastrophic forgetting issues.
@@
Model Editing seeks to edit incorrect or undesirable knowledge encoded in pre-trained models. Some studies initially identify knowledge-related parameters of the existing pre-trained models and then directly edit particular knowledge into these parameters [12, 29, 30, 41, 42]. Other efforts have been explored to store new or correct knowledge in an extra memory, replacing the original predictions with this knowledge during generation [33, 9]. Additionally, meta-learning based methods learn to edit models through meta-learning [2, 32]. However, these approaches are only applicable to modifying specific knowledge. In contrast, our method is independent of specific knowledge: regardless of the type of knowledge contained in the context, it enhances the LLM’s utilization of this knowledge.
@@ 3 Methodology
We focus on tasks involving context-specific knowledge conflicts. The input of these tasks is formulated as $(c,q)$, where $c$ is the context, and $q$ represents the question in completion tasks or the question combined with a suffix consisting of choices in multiple-choice tasks. We propose a novel method that is dedicated to improving the faithfulness of LLMs to the context to address these tasks. The proposed IRCAN methodology is structured into three phases: Initially, we compute the attribution scores of each neuron to assess its influence on context processing. Subsequently, neurons that are responsive to context, termed “context-aware neurons”, are identified. In the final step, we enhance the influence of these detected neurons through a reweighting process, thereby augmenting their impact on the model’s generation. The framework of IRCAN is illustrated in Figure 1.
Figure 1: The diagram of IRCAN. When an LLM faces a knowledge conflict between the context and its inherent knowledge, IRCAN first calculates the attribution score for each neuron to measure its contribution to processing the context. It then identifies context-aware neurons by taking the intersection of neurons with the highest scores. Subsequently, the identified neurons are reweighted so that IRCAN could guide the model to be more aligned with the contextual knowledge, ensuring greater fidelity to the context.
@@ 3.2 Context-Aware Neuron Identification
Based on the calculated neuron attribution scores $\mathrm{Attr}(n_{i}^{l})$, we first retain neurons with scores above the threshold $t$, creating a coarse set of context-aware neurons. Then, for each example, we select the top $z$ neurons with the highest attribution scores as the candidate set. In our experiments, $t$ and $z$ are set to 10% and 20 respectively. Ultimately, we count the number of co-occurrences of neurons in all candidate sets, and we select the top $h$ neurons with the highest number of co-occurrences as context-aware neurons. These identified context-aware neurons are shared across all data instances.
@@ 4.1 Dataset
For the multiple-choice task, we utilized the COSE_KRE and ECARE_KRE datasets [46], which were derived from ECQA [1] and e-CARE [10], respectively. The derivation process involved selecting one of the incorrect answer choices and prompting ChatGPT to generate explanations supporting this incorrect answer. Specifically, the selected incorrect answer was treated as the correct answer, and the explanations generated by ChatGPT were used as the context for the multiple-choice question.
@@ 4.2 Models and Metrics
Although we used accuracy as the primary metric to evaluate our method more comprehensively, we also designed a supplementary metric: stubbornness rate (SR), which measures whether the LLM persistently adheres to its internal memorized knowledge. This metric is defined as the accuracy with which the model’s generation matches the original golden label (for the MemoTrap dataset, i.e., the common ending word of a well-known proverb; for the COSE_KRE and ECARE_KRE dataset, i.e., the original golden option). A lower stubbornness rate indicates that the LLM exhibits a decreased reliance on the knowledge encapsulated within its internal parameters during the generation process, suggesting a greater propensity to incorporate contextual information to a certain extent.
@@
ITI (Probe Weight Direction) 26.01 25.16 31.77 44.78 20.26 43.50 18.34 53.52 23.03 51.17
ITI (Mass Mean Shift) 0.00 0.00 31.34 44.99 0.00 0.00 18.12 53.94 22.60 52.45
@@
ITI (Probe Weight Direction) 31.59 23.57 37.32 17.51 40.75 45.01 50.41 25.37
ITI (Mass Mean Shift) 29.46 23.73 26.35 18.66 38.95 43.04 25.20 19.15
@@
ITI (Probe Weight Direction) 73.04 26.96 49.58 50.42 60.51 39.49 73.42 26.58
ITI (Mass Mean Shift) 73.80 26.20 47.60 52.40 49.58 50.42 71.44 28.56
@@ 4.3 Baselines
To demonstrate the effectiveness of IRCAN, we compared it with the following baselines: Original: which refers to the LLMs without any modification. Prompt engineering based methods: we curated three types of prompts to explicitly instruct LLMs to pay more attention to the knowledge in context on the multiple-choice task. According to the content added in the prompt, we denote these three prompt engineering based methods as Based_on, Based_on_Formatted, and Utilizing_Formatted, respectively. Please refer to Appendix C for the details of these prompts. Inference-Time Intervention (ITI) [22]: this method identifies a direction in the activation space associated with factually correct statements and shifts activations along this direction during inference, thereby enhancing the truthfulness of LLMs. Analogous to its experimental setup, for each sample in each dataset, we concatenated the question/answer together and extracted head activations at the last token to collect a probing dataset. We then used the ITI method to identify the direction and intervened activations along this direction. We implemented two intervention directions in our experiments, i.e., Probe Weight Direction and Mass Mean Shift. Context-aware decoding (CAD) [36]: it adjusts the output probabilities of LLMs to emphasize differences when the context is utilized versus when it is absent. Specifically, this approach reduces the weighting of prior knowledge in favor of more pertinent contextual information. Since this method and our approach work in completely different ways: CAD manipulates the model’s final output probabilities, while IRCAN focuses on identifying and enhancing neurons responsible for processing the context, these two methods can be seamlessly combined without any obstacles. In our experiments, we also report the performance of the combined system, which is referred to as IRCAN-CAD in our experiments. Opin [49]: this method argues that transforming factual questions into questions seeking opinions allows people to pay more attention to the context. Therefore, it uses opinion-based prompts to enable the model to generate context-faithful responses. Since the prompts designed by Opin do not apply to the completion task, we used this method as a baseline only on the multiple-choice datasets. To ensure a fair comparison, we adopted the same method for selecting the generated answer as described in §4.2 for Opin.
@@ 4.4 Main Results
We treat the number of identified context-aware neurons $h$3.2), ranging from 1 to 16, and the enhancement strength $\beta$ for these neurons (§3.3), ranging from 2 to 20, as hyperparameters. All datasets were randomly divided into validation and test sets in a 1:1 ratio. We employed grid search to identify the optimal hyperparameter configuration that maximized performance on the validation set. Subsequently, we report the test set results obtained using the identified optimal hyperparameter combination in Table 1 and Table 2.
@@
The main results on the MemoTrap dataset are shown in Table 1. IRCAN significantly outperforms all the baselines. Intervening along the Mass Mean Shift direction brings different degrees of performance degradation to the vast majority of LLMs. Even more, the interventions cause Gemma-2B and Amber to completely fail to respond normally. Improvements along the Probe Weight Direction are also limited. This suggests that finding a direction relevant to the context and shifting activations along this direction during inference might not be a good way to enhance the attention of an LLM to contextual knowledge. The CAD method shows a slightly better improvement, but there is still a significant gap compared to our IRCAN. Notably, IRCAN achieves remarkable improvements of 129% for LLaMA-2-7B and 136% for LLaMA-3-8B in terms of accuracy when compared to the Original baseline. Such substantial performance improvements, achieved through strengthening merely a few or a dozen neurons, indicate that the neurons identified by our method play a pivotal role in processing context.
@@
It is also notable that, in the Original setting, LLaMA-3-8B achieves the lowest ACC, along with the highest SR. This observation may seem somewhat counterintuitive. We believe it is due to the fact that LLaMA-3-8B, trained on extensive, high-quality multi-source data, has acquired more extensive world knowledge and tends to rely more on its pre-stored intrinsic knowledge when generating responses.
@@
As presented in Table 2, ITI still performs poorly on the multiple-choice task. Moreover, only instructing LLMs to pay more attention to the knowledge in the context is not sufficient to enhance the model’s utilization of contextual knowledge. Furthermore, the prompts with the best performance differ for different datasets and different models. Therefore, the requirement of meticulous prompt engineering undermines their generalizability.
Figure 4: Model performance with different enhanced # neurons $h$.
In contrast, IRCAN significantly enhances the performance of LLMs in resolving knowledge conflicts in the multiple-choice task. This improvement is observed across the COSE_KRE and ECARE_KRE datasets for the majority of the evaluated LLMs. The multiple-choice task poses a greater challenge compared to the MemoTrap dataset, where the context directly provides the required knowledge for generation. In the multiple-choice task, LLMs are required to interpret and reason based on the implicit knowledge in the context to facilitate generation. Consequently, the improvements in performance observed for the two baseline models and our proposed IRCAN are less pronounced on this dataset than those noted on MemoTrap. Nonetheless, IRCAN still achieves a gain in accuracy and a drop in SR for the majority of LLMs engaged in this task, setting new state-of-the-art results.
@@ 5.1 Effect of Enhancement Strength
The relationship between the accuracy of the LLaMA-2-7B model on the MemoTrap dataset and the enhancement strength $\beta$ is depicted in Figure 3, where the number of enhanced neurons is fixed at 14. For further results on various enhancement strengths $\beta$, please refer to Figure 11 in Appendix D. It can be observed that as the enhancement strength for the context-aware neurons increases, model performance gradually improves, highlighting the pivotal role of the neurons identified by our method. Then, consistent with our intuition, performance begins to decline beyond a certain enhancement strength (7 in this scenario). This decline could be due to excessively high enhancements of certain neurons, leading to uncontrollable outputs or a reduction in model capabilities.
@@ 5.2 Effect of the Number of Context-Aware Neurons
The impact of the number of enhanced neurons $h$ on the performance of LLaMA-2-7B on the MemoTrap dataset is shown in Figure 4. We present results where the enhancement multiplier is fixed at 7, with comprehensive results available in Appendix D (see Figure 12). Similarly, observations reveal that as the number of enhanced neurons increases, the model’s accuracy initially improves but subsequently begins to decline, resonating with the results observed with the enhancement strength.
@@ 5.3 Layer Distribution of Context-Aware Neurons
We also analyzed the overlap of identified context-aware neurons across different prompts. Results (Appendix E) show a high degree of overlap, demonstrating IRCAN consistently identifies these neurons regardless of the prompt.
@@ 5.4 Evaluation of General Abilities of LLMs
To investigate whether up-weighting context-aware neurons impairs the model’s general abilities, we conducted evaluations of IRCAN on six widely-used benchmarks. These benchmarks are used in the widely-recognized Open LLM Leaderboard, [^3] https://huggingface.co/spaces/HuggingFaceH4/open_llm_leaderboard including ARC [6], HellaSwag [47], MMLU [15], Winogrande [34], GSM8K [7], and TruthfulQA [24]. We describe the experimental details in Appendix F. The experimental results, as shown in Table 3, reveal that IRCAN rarely impacts the general ability of the LLMs. Surprisingly, in some cases, it even leads to a slight performance improvement. These results suggest that IRCAN can reliably improve the capability of the LLMs in addressing knowledge conflict tasks while maintaining their excellent general capabilities.